Our website has detected that you are using a browser that is not supported.
This may prevent you from accessing certain features.
We strongly recommend you use an alternative browser.
We have fully test this site on both Chrome and Firefox. Which are available for download here
Sensor Info
Sensor Status
Time of last reading:
March 20, 2019, 3:29 p.m.
First Record:
June 1, 2017, noon
Sensor Issues
Issue Jan. 30, 2018 - Jan. 30, 2018
NO sensor replaced
Issue Nov. 18, 2017 - Nov. 19, 2017
Third party server outage. All AQ Mesh and precision sites affected.
Issue Feb. 12, 2018 - Feb. 22, 2018
It has been brought to our attention by Environmental Instruments Ltd (manufacturers of AQMesh) that due to a bug in a server systems update that NO2 and O3 data from all AQMesh pods were effected from 7am on 12th February until 9am on 22nd February 2018. No other channels are affected. We caution you against using NO2 and O3 data for this period until the corrections have been applied. We urge users to re-call this data from Tuesday 27th when all readings will have been re-processed. The Urban Observatory apologises for the inconvenience.
Issue Dec. 4, 2017 - Dec. 4, 2017
NO and NO2 sensors replaced after a fault
Issue Sept. 17, 2018 - Oct. 27, 2018
Between these dates, no data was sent by this pod to AirMonitors on any channel. Rebooted on-site, data came back.
Issue Oct. 11, 2018 - Oct. 17, 2018
During this period, data was intermittent/missing Battery changed and restarted at 11:00 17/10/18
Issue Oct. 24, 2018 - Oct. 25, 2018
All AQ Meshes did not report any data between Weds 24th Oct 06:00 – Thurs 25th Oct 11:15, reason not yet known
Issue Nov. 16, 2018 - Nov. 16, 2018
Approx 15:00 - 16:00 : Changed the sampling frequency for gases for all AQ Meshes to be standard across all pods : P1 (Gas Reading Frequency) - Ten secs (unchanged) P2 (Gas Averaging Frequency) - One Minute (was mostly 15 mins) P3 (Gas Transmission Frequency) - 30 Mins (was One Hour)
Issue Nov. 22, 2018 - Nov. 23, 2018
22/11/18 17:00 AirMonitors carrying out server maintenance at approx 17:00 (UTC) on Thursday November 22nd 23/11/10:00 server back. Possible data loss during this period
Issue Oct. 20, 2018 - Nov. 27, 2018
Missing data between 20/10 and 27/11. A new large sign has been placed directly next to the panel, together with the low winter sun and a faulty battery, this has caused problems keeping charging. 27/11/18 05:30 New panel and battery installed, and the panel has been oriented more towards the open sky (ie ESE, as South-facing is obscured by Northern Goldsmiths)
Issue Dec. 6, 2018 - Dec. 6, 2018
06/12/2018 Major disruption to the 02 Cellphone/Data network today. This will affect many sensors that rely on that network (eg Emotes)
Issue Jan. 20, 2019 - Jan. 21, 2019
20/01/19 17:00 approx Sensors data not being displayed on the portal. This was due to failed video data stuck in the queue and blocking all other data.
Issue Jan. 15, 2019 - Jan. 16, 2019
15/01/19 02:15 - 16/01/19 Data lost from all sensors due to quue failure. This data may eventually get back-filled.
Sensor Notes:
Sensor Maintenance Records:
Installation Data:
Maintenance checks: